OpenStudy (anonymous):

ABC is a triangle, in which 2sinB.cosC+sinC=sinA+sinB , find AC/CB

OpenStudy (anonymous):

@Directrix ?

OpenStudy (anonymous):

@TuringTest ?? @UnkleRhaukus ?? @Hero ?? @ParthKohli ??

OpenStudy (anonymous):

@phi ?? @experimentX ??

OpenStudy (experimentx):

try making use of sine laws

OpenStudy (experimentx):

also you know sin(A+B+C) = 0

OpenStudy (experimentx):

woops!! that was sides ... hold on a bit,

OpenStudy (experimentx):

looks like you can make use of it. http://en.wikipedia.org/wiki/Law_of_sines from there get the value of sin(B)/sin(A) in terms of sin(C) from sin(A+B) = -sin(C) or sin(A+B+C) = 0 ... get again the value of sin(B)/sin(A) in terms of sin(C)

OpenStudy (experimentx):

solve for the value of sin(C) ... and put is there to get the value of that ratio ... from sine laws you have the value of that ratio. Sorry can't do whole stuff .. busy for a while.

OpenStudy (anonymous):

how would we start?

OpenStudy (anonymous):

\[\cos c=?\]

OpenStudy (experimentx):